How do you put [x=y^2+8y+13] into vertex form and how do you graph it?

Answer by stanbon(48569)

How do you put [x=y^2+8y+13] into vertex form and how do you graph it?
Square the y-terms to get:
x=y^2+8y+16+13-16
x=y^2+8y+16-3
x+3 = (y+4)^2
Vertex at (-3,-4)
x-intercept at 13
y-intercepts:
y = [-8+-sqrt(8^2-4*13)]/2
y = [-8+-2sqrt3]/2
y = (-4+sqrt3) or y = -4-sqrt3
-------------
The parabola opens to the right; has vertes, x-intercept, and y-intercepts
as indicated.
===================
Cheers,
Stan H.